OpenStudy (zolock):

5. Find the values of k that make 2x2 + kx + 11 factorable. a=1 or 11 b=11 or 13 c=13 or 23 d=2 or 11

OpenStudy (campbell_st):

for any quadratic \[ax^2 +bx + c \] b is the sum of the factors of a*c so you need to find the factors of 2*11 = 22 so they can be 2 and 11 or 1 and 22... then b = 13 or b = 23 so the quadratic would become \[2x^2 +13x + 11 = (x + 1)(2x + 11)\] or \[2x^2 + 23x + 11 = (2x + 1)(x + 11)\] hope it helps